# Linearly Polarized Light-Induced Anomalous Hall Effect and Topological Phase Transitions in an Altermagnetic Topological Insulator

### 06 Mar 2026 v1

@@ I Introduction
Among various approaches to controlling electronic structures, periodic light irradiation offers a powerful dynamical route known as Floquet engineering, enabling access to diverse light-induced phenomena[55, 100, 107, 38, 113, 57, 28, 86, 19, 31, 64, 74, 69, 18, 63, 99, 7, 70, 27, 94, 5, 87, 6, 52, 68, 62, 14, 20, 49, 116, 110, 73, 41, 67, 106, 42, 72, 101, 108, 117, 8, 9]. Recently, light-induced manipulation of electronic and magnetic properties in AMs has attracted growing interest, such as light-induced odd-parity magnetism[118, 37, 56, 53], light-induced higher-order SOC[26], Floquet engineering spin triplet states[24], circularly polarized light (CPL) induced quantum Anomalous Hall Effect (QAHE) [119] and so on[25, 80, 98, 77, 47, 109]. In contrast to previous proposals where light-induced anomalous Hall effects (AHE) predominantly rely on CPL to break $\mathcal{T}$ symmetry and parity-time-reversal ($\mathcal{PT}$) symmetry[100, 73, 41, 67, 106, 42, 72, 101, 108, 117], linearly polarized light (LPL) is typically ineffective in nonmagnetic and conventional antiferromagnetic materials. AMs circumvent this limitation owing to their intrinsically broken $\mathcal{T}$ and $\mathcal{PT}$ symmetries. Moreover, AMs themselves can host nontrivial band topology, and AM topological insulator phases have been explored [95, 82, 61, 2, 81, 78, 23, 51, 22, 12, 111, 29, 30]. Despite these advances, the interplay between LPL driving and the underlying topology and magnetism of AMs remains largely unexplored.
@@
In this work, we employ a four-band model on a two-dimensional square lattice with $d$-wave spin splitting, and derive the effective Hamiltonian under LPL irradiation within Floquet theory. We calculate and analyze the band structures, anomalous Hall conductivity (AHC), and topological phase transitions induced by LPL in out-of-plane $d$-wave AMs. We further compare these results with out-of-plane AFMs under LPL irradiation. We find that LPL breaks the lattice symmetry of the AMs and lifts the band degeneracy along high-symmetry lines. Furthermore, we demonstrate that although LPL does not break $\mathcal{T}$ or $\mathcal{PT}$ symmetries, it can nevertheless induce the AHE in AMs. Intriguingly, we reveal that under LPL irradiation, the AM topological insulator can be tuned to a spin-polarized Chern insulator phase as the light intensity increases, thereby exhibiting the QAHE with a single spin-polarized. Such phenomena are absent in AFMs under identical irradiation conditions. Our work not only uncovers further distinctions between AMs and AFMs but also proposes a method for experimentally distinguishing between them.
Figure 1: Schematic illustrations of the optical setup and the magnetic lattice models. (a) Top view of the linearly polarized light configuration, where $\theta$ denotes the polarization angle relative to the $x$ axis. (b) Lattice model of the two-dimensional $d$-wave altermagnet. (c) Lattice model of the two-dimensional antiferromagnet. In (b) and (c), red and blue solid circles represent atoms with spin-up and spin-down moments, respectively, while gray circles represent non-magnetic atoms.
@@ II Model and Method
To investigate the control of AM via Floquet engineering with LPL, we first employ an out-of-plane $d$-wave AM model and subsequently utilize Floquet theory to derive the effective model under LPL irradiation.
@@ II.1 The d-wave AM tight-binding model
where $H_{\downarrow}(k_{x},k_{y})=H_{\uparrow}(k_{y},k_{x})$, $H_{\uparrow}(\mathbf{k})=\sum_{i}{d_{i}}^{\uparrow}(\mathbf{k})\tau_{i}$ with $i=x,y,z$, $\tau_{i}$ denote the Pauli matrices acting on the orbital subspace, $d_{x}^{\uparrow}(\mathbf{k})=-v\sin k_{x}$, $d_{y}^{\uparrow}(\mathbf{k})=-vt_{a}\sin k_{y}$, $d_{z}^{\uparrow}(\mathbf{k})=m+b(\cos k_{x}+t^{2}_{a}\cos k_{y})$. Here, $m$ corresponds to the on-site atomic potential, which depends on the local magnetic moment, $v$ describes the hopping amplitude between different orbitals. For simplicity, we set $v=1$. The constant $t_{a}$ introduces a symmetry breaking between the $x$ and $y$ directions within each spin block, rendering the Hamiltonian characteristic of a $d$-wave AM when $t_{a}\neq 1$. The spin-up and spin-down channels are related by $C_{4z}\mathcal{T}$ or $\mathcal{M}_{xy}$. Here $C_{4z}$ is the fourfold rotation about the $z$ axis, and $\mathcal{M}_{xy}$ is the mirror symmetry. When $t_{a}=1$, the system restores $\mathcal{PT}$ symmetry and reduces to a conventional AFM as shown in Fig. 1(c). The Hamiltonian in Eq. (1) describes a quantum spin Hall (QSH) insulator when $|m|<|b|(1+t_{a}^{2})$, otherwise, it is topologically trivial.
@@
Figure 2: Calculated band structures and Fermi surfaces of altermagnets (AMs) and antiferromagnets (AFMs). Band structures (first and third columns) and Fermi surfaces at $E_{F}=2v$ (second and fourth columns) are shown for the AM [panels (a), (b), (e), and (f)] and AFM [panels (c), (d), (g), and (h)]. The rows correspond to different light intensity. (a)-(d) the equilibrium state without light irradiation ($A_{0}=0$); (e)-(h) under irradiation by linearly polarized light (LPL) with light intensity at $A_{0}=1.0$ and polarization direction at $\theta=0$. Red solid lines and blue dashed lines represent the spin-up and spin-down bands, respectively. The parameters are set as: $m=4.2v$, b=vb=-v and $t_{a}=\sqrt{3}$ for the AM; and $m=4.2v$, b=2vb=-2v, and $t_{a}=1$ for the AFM.
@@ II.2 Effective Hamiltonian with Floquet theory
For a periodically driven system, the Hamiltonian satisfies $H(\mathbf{k},t)=H(\mathbf{k},t+T)$, where $T=\frac{2\pi}{\omega}$ is the drive period. According to the Floquet theory, the wavefunction of a periodically driven system satisfies $\psi_{\alpha}(\mathbf{k},t)=e^{-i\epsilon_{a}t}u_{\alpha}(\mathbf{k},t)$, in which $\epsilon_{\alpha}$ is the quasienergy and $u(\mathbf{k},t)$ is periodic, namely, $u_{\alpha}(\mathbf{k},t)=u_{\alpha}(\mathbf{k},t+T)$. The periodic function $u_{\alpha}(\mathbf{k},t)$ satisfies the Schrödinger equation $H(\mathbf{k},t)-i\partial_{t}]u_{\alpha}(\mathbf{k},t)=\epsilon_{\alpha}u_{\alpha}(\mathbf{k},t)$. Due to the periodicity, $u_{\alpha}(\mathbf{k},t)$ and $H(\mathbf{k},t)$ are expanded as $u_{\alpha}(\mathbf{k},t)=\sum_{n}u_{n\alpha}(\mathbf{k})e^{in\omega t}$ and $H(\mathbf{k},t)=\sum_{n}H_{n}(\mathbf{k})e^{in\omega t}$, $u_{n\alpha}(\mathbf{k})$ and $H_{n}(\mathbf{k})$ are the nth Fourier component of $u_{\alpha}(\mathbf{k},t)$ and $H(\mathbf{k},t)$. The quasienergy $\epsilon_{\mathbf{k}\alpha}$ satisfies the following equation:
@@
$\bm{A}(t)=A_{0}(\cos{\theta}\cos(\omega t),\sin{\theta}\cos(\omega t)),$ (5)
where $A_{0}$ is the amplitude of the LPL, $\theta$ denotes the angle between the polarization direction and the x-axis, as illustrated in Fig. 1(a). The time-dependent Hamiltonian is obtained by making Peierls substitution $\mathbf{k}\rightarrow\mathbf{k}+e\bm{A}(t)/\hbar$. It is noted that the vector potential of LPL satisfies $\bm{A}(t)=\bm{A}(-t+\tau)$, where $\tau$ is the fractional time translation, which results in the Fourier component $H_{n}(\mathbf{k})=e^{-in\omega\tau}H_{-n}(\mathbf{k})$, the term $\sum_{n\geq 1}\frac{[H_{n},H_{-n}]}{n\omega}$ vanishes, yielding $H_{\text{eff}}(\mathbf{k})\approx H_{0}(\mathbf{k})$. By applying the Jacobi-Anger expansion, the effective model Hamiltonian for the $d$-wave AM driven by LPL at the high-frequency limit is derived as:
@@ II.3 Numerical Calculation of the AHC
where $n$ is band index, $\hat{v}_{x}$$\frac{\partial H_{(}\mathbf{k})}{\partial k_{x}}$ and $\hat{v}_{y}$$\frac{\partial H(\mathbf{k})}{\partial k_{y}}$ are velocity operators, $f_{n}(\mathbf{k})$ is the Fermi-Dirac distribution function. $\Omega_{n}(\mathbf{k})$ denotes the Berry curvature, $\varepsilon_{n\mathbf{k}}$ and $\lvert\phi_{n\mathbf{k}}\rangle$ are the energy eigenvalue and eigenstate of the Floquet effective Hamiltonian, respectively.
@@ III.1 Modulation of electronic structures by LPL
We first examine the LPL-induced modulation of the symmetry and electronic structure in AFM and AM. When $t_{a}=1$, the system preserves $\mathcal{PT}$ symmetry, representing a conventional AFM. The system exhibits Kramers degeneracy throughout the entire Brillouin zone (BZ), where the spin-up and spin-down energy bands are completely degenerate[See Figs. 2(c) and (d)]. The spin-up and spin-down bands in the AFM remain degenerate under LPL irradiation due to the preserved $\mathcal{PT}$ symmetry, as shown in Figs. 2(g) and (h).
When $t_{a}\neq 1$, the $\mathcal{PT}$ symmetry is naturally broken, and the system transitions into a d-wave AM where the spin-up and spin-down bands are related by the $C_{4z}\mathcal{T}$ or $\mathcal{M}_{xy}$ symmetries. Due to the $C_{4z}\mathcal{T}$ and $\mathcal{M}_{xy}$ symmetries, the spin-up and spin-down bands are degenerate along the path $\Gamma$-M. However, LPL with arbitrary polarization directions does not preserve the $C_{4z}\mathcal{T}$ and $\mathcal{M}_{xy}$ symmetries, lifting the degeneracy along the $\Gamma$-M path, as shown in Fig. 2(e). Furthermore, the spin-up and spin-down bands are no longer connected by any symmetry operation. This leads to the transition from an AM to a compensated ferrimagnet.
Figure 3: Light-induced anomalous Hall effect (AHE). (a) The anomalous Hall conductivity (AHC) as a function of the Fermi energy for $A_{0}=0$ (black line), $A_{0}=0.3$ (blue line), $A_{0}=0.6$ (red line) and $A_{0}=0.9$ (green line), with $\theta=0$. (b) Comparison between the AHC calculated using the high-frequency approximation (red line) and accounting for the occupation of Floquet states (blue line) at $\omega=10$, with $A_{0}=0.6$ and $\theta=0$. (c) The AHC as a function of $\theta$ at a fixed Fermi energy $E_{F}=2v$. (d) Schematic polar plot sign of $\sigma_{xy}$ for AMs. Red (blue) regions indicate positive (negative). The parameters are $m=4.2v$, b=vb=-v, and $t_{a}=\sqrt{3}$.
@@ III.2 LPL-induced AHE
In $\mathcal{PT}$-symmetric AFM, $\mathcal{PT}$ symmetry strictly constrains the Berry curvature to zero, so there is no AHE in $\mathcal{PT}$-symmetric AFM. LPL cannot break $\mathcal{PT}$ symmetry, and therefore LPL cannot induce an AHE in AFM. In out-of-plane AM, the spin-up and spin-down electrons are connected by $C_{4z}\mathcal{T}$ or $\mathcal{M}_{xy}$ symmetries enforcing a vanishing AHE. However, LPL can break these symmetries, thereby inducing an AHE. To evaluate the AHE induced by LPL in AM, two methods were employed to determine the AHC, including a direct computation of the Berry curvature integral within the high-frequency approximation and an alternative approach based on Floquet state occupations.
In Fig. 3(a), we calculate the AHC in the AM as a function of the Fermi level under various LPL intensities within the high-frequency approximation. It is evident that for $A_{0}=0$, the AHC of the AM vanishes. However, as the light intensity increases, the $C_{4z}\mathcal{T}$ and $\mathcal{M}_{xy}$ symmetries are broken by the LPL, thereby inducing a finite AHC. The distinct anomalous Hall responses of AFM and AM under LPL can serve as an experimental signature to distinguish between them. To achieve a more accurate calculation of the AHC, the occupations of Floquet states must be taken into account. Following Eq. (11), we calculated the AHC at $\omega=10$. A comparison between these results and those derived from the high-frequency effective Hamiltonian is presented in Fig. 3(b). It is evident that at this frequency, the results from both methods are in excellent agreement, thereby validating the effectiveness of the high-frequency approximation.
@@
Figure 4: Light-induced evolution of the spin-resolved band gap and the associated topological phase transitions under linearly polarized light (LPL). The evolution of the spin-up and spin-down band gaps under LPL irradiation for (a) a trivial altermagnet (AM), (b) a trivial antiferromagnet (AFM), (c) quantum spin Hall (QSH) AM, (d) QSH AFM. The red solid lines and blue dashed lines represent the band gaps for the spin-up and spin-down bands, respectively. Yellow, light blue, and white regions denote the QSH, Chern insulator (C=1C=-1), and topologically trivial phases, respectively. The parameters for a trivial AM are $m=4.2v,b=-v,t_{a}=\sqrt{3}$, for a trivial AFM are $m=4.2v,b=-2v,t_{a}=1$, for a AM QSH insulator are $m=3.8v,b=-v,t_{a}=\sqrt{3}$, for a AFM QSH insulator are $m=3.8v,b=-2v,t_{a}=1$. The polarization direction $\theta=0$.
Figure 5: Phase diagrams under linearly polarized light irradiation as a function of the light intensity $A_{0}$ and the polarization direction $\theta$. (a) The phase diagram for the $d$-wave quantum spin Hall (QSH) altermagnets, the parameters are $m=3.8v,b=-1v,t_{a}=\sqrt{3}$. (b) The phase diagram for the QSH antiferromagnet, the parameters are $m=3.8v,b=-2v,t_{a}=1$.

### 13 Mar 2026 v2

@@ I Introduction
Among various approaches to controlling electronic structures, periodic light irradiation offers a powerful dynamical route known as Floquet engineering, enabling access to diverse light-induced phenomena[55, 101, 108, 38, 116, 57, 28, 87, 19, 31, 64, 75, 69, 18, 63, 100, 7, 71, 27, 95, 5, 88, 6, 52, 68, 62, 14, 20, 49, 119, 113, 74, 41, 67, 107, 42, 73, 102, 109, 120, 8, 9, 70]. Recently, light-induced manipulation of electronic and magnetic properties in AMs has attracted growing interest, such as light-induced odd-parity magnetism[121, 37, 56, 53], light-induced higher-order SOC[26], Floquet engineering spin triplet states[24], circularly polarized light (CPL)-induced quantum anomalous Hall (QAH) Effect [122] and so on[25, 81, 99, 78, 47, 112, 111, 110]. In contrast to previous proposals where light-induced anomalous Hall effects (AHE) predominantly rely on CPL to break $\mathcal{T}$ symmetry and parity-time-reversal ($\mathcal{PT}$) symmetry[101, 74, 41, 67, 107, 42, 73, 102, 109, 120], linearly polarized light (LPL) is typically ineffective in nonmagnetic and conventional antiferromagnetic materials. AMs circumvent this limitation owing to their intrinsically broken $\mathcal{T}$ and $\mathcal{PT}$ symmetries. Moreover, AMs themselves can host nontrivial band topology, and AM topological insulator phases have been explored [96, 83, 61, 2, 82, 79, 23, 51, 22, 12, 114, 29, 30]. Despite these advances, the interplay between LPL driving and the underlying topology and magnetism of AMs remains largely unexplored.
@@
In this work, we employ a four-band model on a two-dimensional square lattice with $d$-wave spin splitting, and derive the effective Hamiltonian under LPL irradiation within Floquet theory. We calculate and analyze the band structures, anomalous Hall conductivity (AHC), and topological phase transitions induced by LPL in out-of-plane AFMs and $d$-wave AMs. We reveal that LPL cannot break the $\mathcal{PT}$ symmetry connecting the spin-up and spin-down sublattices [Fig. 1(a)], and therefore cannot lift the spin degeneracy. As the intensity of the LPL increases, it drives the AFM QSH insulator into a trivial insulator, as schematically illustrated in Fig. 1(b). In contrast, we find that in the AM, where the spin-up and spin-down sublattices are connected by $C_{4z}\mathcal{T}$ symmetry, LPL can break this symmetry[Fig. 1(c)]. Consequently, the spin-up and spin-down bands do not close and reopen simultaneously as the LPL intensity increases. This allows the AM QSH insulator to undergo a transition to a spin-polarized Chern insulator under LPL irradiation, as shown in Fig. 1(d). Furthermore, we observe that the AHE becomes anisotropic with respect to the polarization direction of the LPL, enabling sign reversal upon polarization rotation. Our work not only provides an experimentally feasible means to distinguish between AMs and AFMs but also reveals the different topological phase transitions in AFM and AM topological insulators under LPL irradiation.
Figure 1: Schematic illustrations of linearly polarized light (LPL) control on the altermagnet (AM) and antiferromagnet (AFM). (a) In conventional AFM, the spin-up and spin-down sublattices are connected by $\mathcal{PT}$ symmetry, and incident LPL does not break $\mathcal{PT}$ symmetry. (b) As the light intensity increases, the two-dimensional AFM quantum spin Hall insulator transitions into a trivial insulator. The black lines indicate that the energy bands for spin-up and spin-down are degenerate, and the dashed lines represent band inversion. (c) In the AM, spin-up and spin-down sublattices are connected by $C_{4z}\mathcal{T}$ symmetry, and incident LPL breaks this symmetry. (d) As the light intensity increases, the two-dimensional AM quantum spin Hall insulator first transitions into a spin-polarized Chern insulator and then into a trivial insulator. The red and blue lines represent the energy bands for spin-up and spin-down, respectively.
@@ II Model and Method
To investigate the manipulation of AM via Floquet engineering with LPL, we first employ an out-of-plane $d$-wave AM tight-binding model and subsequently utilize Floquet theory to derive the effective model under LPL irradiation.
@@ II.1 The d-wave AM tight-binding model
where $H_{\downarrow}(k_{x},k_{y})=H_{\uparrow}(k_{y},k_{x})$, $H_{\uparrow}(\mathbf{k})=\sum_{i}{d_{i}}^{\uparrow}(\mathbf{k})\tau_{i}$ with $i=x,y,z$, $\tau_{i}$ denote the Pauli matrices acting on the orbital subspace, $d_{x}^{\uparrow}(\mathbf{k})=-v\sin k_{x}$, $d_{y}^{\uparrow}(\mathbf{k})=-vt_{a}\sin k_{y}$, $d_{z}^{\uparrow}(\mathbf{k})=m+b(\cos k_{x}+t^{2}_{a}\cos k_{y})$. Here, $m$ corresponds to the on-site atomic potential, which depends on the local magnetic moment, $v$ describes the hopping amplitude between different orbitals. For simplicity, we set $v=1$. The constant $t_{a}$ introduces a symmetry breaking between the $x$ and $y$ directions within each spin block, rendering the Hamiltonian characteristic of a $d$-wave AM when $t_{a}\neq 1$. The spin-up and spin-down channels are related by $C_{4z}\mathcal{T}$ or $\mathcal{M}_{xy}$. Here $C_{4z}$ is the fourfold rotation about the $z$ axis, and $\mathcal{M}_{xy}$ is the mirror symmetry. When $t_{a}=1$, the system restores $\mathcal{PT}$ symmetry and reduces to a conventional AFM. The Hamiltonian in Eq. (1) describes a QSH insulator when $|m|<|b|(1+t_{a}^{2})$, otherwise, it is topologically trivial.
@@
Figure 2: (a-c) The band structures and the Fermi surfaces in the equilibrium state. (a), (c) The electronic structures for the altermagnet (AM) (a) and the antiferromagnet (AFM) (c). (b) The corresponding Fermi surfaces at $E_{F}=2v$ for the AM. (d-f) The band structures and the Fermi surfaces under the linearly polarized light irradiation with $A_{0}=1$, $\theta=0$. (d), (f) The electronic structures for the AM (d) and the AFM (f). (e) The corresponding Fermi surfaces at $E_{F}=2v$ for the AM. Red solid and blue dashed lines denote the spin-up and spin-down bands, respectively. The parameters are $m=4.2v$, b=vb=-v, and $t_{a}=\sqrt{3}$ for the AM, and $m=4.2v$, b=2vb=-2v, and $t_{a}=1$ for the AFM.
@@ II.2 Effective Hamiltonian with Floquet theory
For a periodically driven system, the Hamiltonian satisfies $H(\mathbf{k},t)=H(\mathbf{k},t+T)$, where $T=\frac{2\pi}{\omega}$ is the drive period. According to the Floquet theory, the wavefunction of a periodically driven system satisfies $\psi_{\alpha}(\mathbf{k},t)=e^{-i\epsilon_{\mathbf{k}\alpha}t}u_{\alpha}(\mathbf{k},t)$, in which $\epsilon_{\mathbf{k}\alpha}$ is the quasienergy and $u(\mathbf{k},t)$ is periodic, namely, $u_{\alpha}(\mathbf{k},t)=u_{\alpha}(\mathbf{k},t+T)$. The periodic function $u_{\alpha}(\mathbf{k},t)$ satisfies the Schrödinger equation $H(\mathbf{k},t)-i\partial_{t}]u_{\alpha}(\mathbf{k},t)=\epsilon_{\mathbf{k}\alpha}u_{\alpha}(\mathbf{k},t)$. Due to the periodicity, $u_{\alpha}(\mathbf{k},t)$ and $H(\mathbf{k},t)$ are expanded as $u_{\alpha}(\mathbf{k},t)=\sum_{n}u_{n\alpha}(\mathbf{k})e^{in\omega t}$ and $H(\mathbf{k},t)=\sum_{n}H_{n}(\mathbf{k})e^{in\omega t}$, $u_{n\alpha}(\mathbf{k})$ and $H_{n}(\mathbf{k})$ are the nth Fourier component of $u_{\alpha}(\mathbf{k},t)$ and $H(\mathbf{k},t)$. The quasienergy $\epsilon_{\mathbf{k}\alpha}$ satisfies the following equation:
@@
$\bm{A}(t)=A_{0}[\cos{\theta}\cos(\omega t),\sin{\theta}\cos(\omega t)],$ (5)
where $A_{0}$ is the amplitude of the LPL, $\theta$ denotes the angle between the polarization direction and the x-axis. The time-dependent Hamiltonian is obtained by making Peierls substitution $\mathbf{k}\rightarrow\mathbf{k}+e\bm{A}(t)/\hbar$. It is noted that the vector potential of LPL satisfies $\bm{A}(t)=\bm{A}(-t+\tau)$, where $\tau$ is the fractional time translation, which results in the Fourier component $H_{n}(\mathbf{k})=e^{-in\omega\tau}H_{-n}(\mathbf{k})$, the term $\sum_{n\geq 1}\frac{[H_{n},H_{-n}]}{n\omega}$ vanishes, yielding $H_{\text{eff}}(\mathbf{k})\approx H_{0}(\mathbf{k})$. By applying the Jacobi-Anger expansion, the effective model Hamiltonian for the $d$-wave AM driven by LPL at the high-frequency limit is derived as:
@@ II.3 Numerical Calculation of the AHC
where $n$ is band index, $\hat{v}_{x}$$\frac{\partial H(\mathbf{k})}{\partial k_{x}}$ and $\hat{v}_{y}$$\frac{\partial H(\mathbf{k})}{\partial k_{y}}$ are velocity operators, $f_{n}(\mathbf{k})$ is the Fermi-Dirac distribution function. $\Omega_{n}(\mathbf{k})$ denotes the Berry curvature, $\varepsilon_{n\mathbf{k}}$ and $\lvert\phi_{n\mathbf{k}}\rangle$ are the energy eigenvalue and eigenstate of the Floquet effective Hamiltonian, respectively.
@@ III.1 Modulation of electronic structures by LPL
We first examine the LPL-induced modulation of the symmetry and electronic structure in AFM and AM. When $t_{a}=1$, the system preserves $\mathcal{PT}$ symmetry, representing a conventional AFM. The system exhibits Kramers degeneracy throughout the entire Brillouin zone (BZ), where the spin-up and spin-down energy bands are completely degenerate[See Fig. 2(c)]. The spin-up and spin-down bands in the AFM remain degenerate under LPL irradiation due to the preserved $\mathcal{PT}$ symmetry, as shown in Fig. 2(f).
When $t_{a}\neq 1$, the $\mathcal{PT}$ symmetry is naturally broken, and the system transitions into a d-wave AM where the spin-up and spin-down bands are related by the $C_{4z}\mathcal{T}$ or $\mathcal{M}_{xy}$ symmetries. Due to the $C_{4z}\mathcal{T}$ and $\mathcal{M}_{xy}$ symmetries, the spin-up and spin-down bands are degenerate along the path $\Gamma$-M. However, LPL with arbitrary polarization directions does not preserve the $C_{4z}\mathcal{T}$ and $\mathcal{M}_{xy}$ symmetries, lifting the degeneracy along the $\Gamma$-M path, as shown in Fig. 2(d). Furthermore, the spin-up and spin-down bands are no longer connected by any symmetry operation. This leads to the transition from an AM to a compensated ferrimagnet.
Figure 3: Light-induced anomalous Hall effect (AHE). (a) The anomalous Hall conductivity (AHC) as a function of the Fermi energy for $A_{0}=0$ (black line), $A_{0}=0.3$ (blue line), $A_{0}=0.6$ (red line) and $A_{0}=0.9$ (green line), with $\theta=0$. (b) Comparison between the AHC calculated using the high-frequency approximation (red line) and accounting for the occupation of Floquet states (blue line) at $\omega=10$, with $A_{0}=0.6$ and $\theta=0$. (c) The AHC as a function of $\theta$ at a fixed Fermi energy $E_{F}=2v$. (d) Schematic polar plot of the sign of $\sigma_{xy}$ for AMs. The red (blue) regions indicate positive (negative). The parameters are $m=4.2v$, b=vb=-v, and $t_{a}=\sqrt{3}$.
@@ III.2 LPL-induced AHE
In the $\mathcal{PT}$-symmetric AFM, $\mathcal{PT}$ symmetry strictly constrains the Berry curvature to zero, so there is no AHE in the $\mathcal{PT}$-symmetric AFM. LPL cannot break $\mathcal{PT}$ symmetry, and therefore LPL cannot induce an AHE in AFM. In the $d$-wave AM, the spin-up and spin-down electrons are connected by $C_{4z}\mathcal{T}$ or $\mathcal{M}_{xy}$ symmetries enforcing a vanishing AHE. However, LPL can break these symmetries, thereby inducing an AHE. To evaluate the AHE induced by LPL in AM, two methods were employed to determine the AHC, including a direct computation of the Berry curvature integral within the high-frequency approximation and an alternative approach based on Floquet state occupations.
In Fig. 3(a), we calculate the AHC in the AM as a function of the Fermi level under various LPL intensities within the high-frequency approximation. It is evident that for $A_{0}=0$, the AHC of the AM vanishes. However, as the light intensity increases, the $C_{4z}\mathcal{T}$ and $\mathcal{M}_{xy}$ symmetries are broken by the LPL, thereby inducing a finite AHC. The distinct anomalous Hall responses of AFM and AM under LPL can serve as an experimental signature to distinguish between them. To achieve a more accurate calculation of the AHC, the occupations of Floquet states must be taken into account. Following the Eq. (11), we calculated the AHC at $\omega=10$. A comparison between these results and those derived from the high-frequency effective Hamiltonian is presented in Fig. 3(b). It is evident that at this frequency, the results from both methods are in excellent agreement, thereby validating the effectiveness of the high-frequency approximation.
@@
Figure 4: Light-induced evolution of the spin-resolved band gap and the associated topological phase transitions under linearly polarized light (LPL). The evolution of the spin-up and spin-down band gaps under LPL irradiation for (a) a trivial altermagnet (AM), (b) a trivial antiferromagnet (AFM), (c) quantum spin Hall (QSH) AM, (d) QSH AFM. The red solid lines and blue dashed lines represent the band gaps for the spin-up and spin-down bands, respectively. Yellow, light blue, and white regions denote the QSH, Chern insulator (C=1C=-1), and topologically trivial phases, respectively. The parameters for a trivial AM are $m=4.2v,b=-v,t_{a}=\sqrt{3}$, for a trivial AFM are $m=4.2v,b=-2v,t_{a}=1$, for an AM QSH insulator are $m=3.8v,b=-v,t_{a}=\sqrt{3}$, for an AFM QSH insulator are $m=3.8v,b=-2v,t_{a}=1$. The polarization direction $\theta=0$.
Figure 5: Phase diagrams under linearly polarized light irradiation as a function of the light intensity $A_{0}$ and the polarization direction $\theta$. (a) The phase diagram for the $d$-wave altermagnetic (AM) quantum spin Hall (QSH) insulator , the parameters are $m=3.8v,b=-1v,t_{a}=\sqrt{3}$. (b) The phase diagram for the antiferromagnetic (AFM) QSH insulator, the parameters are $m=3.8v,b=-2v,t_{a}=1$.